OpenStudy (anonymous):

3x+y=16 x+4y=20 please show steps thank you

OpenStudy (anonymous):

i got the other now struggleing with another and getting thesteps to work sorry to bother

OpenStudy (jackellyn):

Solve for y 3x+y =16 y=16 -3x Substitute y into second equation x +4(16-3x)=20 x +64 -3x =20 -2x +64 =20 -2x = -44 x = 22 Substitute x into first equation 3(22) + y =16 66 +y =16 y = -50

OpenStudy (anonymous):

your an angel thank you ill need to walk threw the ceck ste with ou however ill have to do t when i get back from my appoinment.

OpenStudy (jackellyn):

missed a part, this is right. x +4(16 -3x) =20 x + 64 -12x =20 -11x +64 =20 -11x = -44 x = 4 3(4) +y =16 12 + y =16 y = 4

OpenStudy (anonymous):

wha is the ste to check it hon